OpenStudy (anonymous):

Four objects, a thin right-cylindrical shell of radius R, a solid right-cylinder of radius R, a thin spherical shell of radius R, and a solid sphere of radius R, each have a mass M. Rank the moments of inertia of each of these object in increasing order from smallest (#1) to largest (#4). The two cylinders are rotated about their cylindrical axes and the two spheres are rotated about any diameter. I said Solid sphere, spherical shell, solid cylinder, cylindrical shell...but apparently that is wrong. I thought you were supposed to compare the equations that all look something like MR^2

OpenStudy (anonymous):

Take a look here: http://en.wikipedia.org/wiki/List_of_moments_of_inertia

OpenStudy (anonymous):

Thats the list i used...I got the result I said in the question

OpenStudy (anonymous):

Let's take a closer look. Hollow Cylinder: \(I=mr^2\) Solid Cylinder: \(I = (mr^2)/2\) Hollow Sphere: \(I=(2mr^2)/3\) Solid Sphere: \(I=(2mr^2)/5\) \(mr^2\) is the same in each, let's compare the fractions. Hollow Cylinder: \(I=1\) Solid Cylinder: \(I=1/2\) Hollow Sphere: \(I=2/3\) Solid Sphere: \(I=2/5\) This leads me to the same answer that you've got. SS, SC, HS, HC

OpenStudy (anonymous):

Actually. Double check your answer. 2/3>1/2

OpenStudy (anonymous):

ah herp derp

OpenStudy (anonymous):

thanks a bunch

OpenStudy (anonymous):

Yep